OpenStudy (anonymous):

The half-life of a certain radioactive material is 76 hours. An initial amount of the material has a mass of 794 kg. Write an exponential function that models the decay of this material.

OpenStudy (anonymous):

794(1/2)^(t/76)

OpenStudy (shamim):

half life T=76 hrs initial amount of radioactive material N0=794 kg

OpenStudy (shamim):

we know radioactive constant lamda=1/T=1/76

OpenStudy (shamim):

N=N0 e^(-lamda *t) N=76 e^(-1/76 *t)